OpenStudy (anonymous):

Solve for y: 2/3 y + 2 = 8 y = -9 y = 11 y = -11 y = 9

sam (.sam.):

\[2/3 y + 2 = 8\] \[\frac{2}{3y+2}=8\]????

OpenStudy (anonymous):

No, It's :